How much will you have to save each month to have $6,000 in two years if the interest rate is 18% compounded monthly?

$398.01

$209.54

$476.59

$299.54
Answer #1

$209.54

Saving each month = Future Value of cash flows / Future Value of annuity of 1
= $       6,000 / 28.63352
= $     209.54
Working:
Future Value of annuity of 1 = (((1+i)^n)-1)/i Where,
= (((1+0.015)^24)-1)/0.015 i = 18%/12 = 0.015
= 28.633521 n = 2*12 = 24